# Evaluating the Impact of Inter-cluster Communications in Edge Computing

### v2

@@ VI Conclusion
In more general terms, we can say that for inter-cluster communication, creating network layer tunnels like Submariner seems to provide higher efficiency for small payloads but suffers when dealing with large amounts of data as soon as significant delay, variance and packet loss is experienced on the connecting link. Solutions with the ability of proxying TCP connections are at advantage here, delivering lower overall response times. Furthermore, use cases where data has to be sent to another service with significantly sized replies seem to suffer when using Skupper or Submariner in edge networks with suboptimal inter-node connections.

### v3

@@ VI Conclusion
In more general terms, we can say that for inter-cluster communication, creating network layer tunnels like Submariner seems to provide higher efficiency for small payloads but suffers when dealing with large amounts of data as soon as significant delay, variance and packet loss is experienced on the connecting link. Solutions with the ability of proxying TCP connections are at advantage here, delivering lower overall response times. Furthermore, use cases where data has to be sent to another service with significantly sized replies seem to suffer when using Skupper or Submariner in edge networks with suboptimal inter-node connections, showing that using HTTP for traffic transport leads to better performance.